A resonant circuit in a radio receiver is tuned to a certain station when the inductor has a value of 0.119 mH and the capacitor has a value of 32.4 pF. Calculate the frequency of the radio station. 2.563E-6Hz. Submit Answer Incorrect. Tries 2/20 Previous Tries What is the wavelength of the radiowaves emited by the radio station? Submit Answer Tries 0/20 Post Discussion Send Feedback
Answer #1

Frequency of a LC resonant circuit is given by

\small f = \frac{1}{2\pi \sqrt{LC}}

here

L = 0.119 mH = 0.119 *10-3 H

C = 32.4 pF = 32.4 * 10-12 m

\small f =\frac{1}{2\pi *0.621* 10^{-7}} = \frac{1}{3.89988*10^{-7}} = 2.564 * 10^{6} Hz  

   \small Wavelength = \lambda = c/f = \frac{3* 10^8}{2.564 *10^6} = 117 m